How to convert 3.3 electron volts to jouls to find its thrishold frewquency

Grade:11

1 Answers

Arun
4 years ago
Dear student
 
3.3 eV = 3.3 * 1.6 * 10^-19 = h (nu)
where nu is frequency
 
52.8 * 10^-19 /6.6 *10^-34 = 8 * 10^15 Hz
 
Hope it helps